Write a structured algorithm to prompt the user to enter the length of a square. Calculate the output and area of the square
1
Expert's answer
2020-11-16T23:34:29-0500

Let p - length of a square (perimeter), so it's side a = p/4, and it's square s = p^2/16:

p = input()
a = p/4
s = a^2
print(s)